# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v2

@@ 3.1 Creation Process
Massive-STEPS is derived from the two subsets of STD [29], incorporating check-ins from both the 2013 and 2018 subsets. We utilize two additional components from STD: (1) the cities metadata file, which provides the latitude and longitude of administrative regions (e.g., towns, suburbs) along with their corresponding country codes obtained from GeoNames; and (2) the POI category mapping, which links each Foursquare Category ID to its descriptive name (e.g., "Restaurant"). Based on this metadata, each POI is thus associated with several attributes: Foursquare Place ID, Foursquare Category ID, category name, latitude/longitude of the administrative region, the administrative region name, and the country code. For anonymization purposes and model training compatibility, we apply ordinal encoding to the Place IDs and Category IDs, assigning each a unique integer index.
Figure 2: World map highlighting the cities included in the Massive-STEPS dataset.
3.1.1 Preprocessing
@@
Most POI recommendation models operate on sequences of check-ins, commonly referred to as trajectories. The model is tasked with predicting the next POIs a user is likely to visit, given the current trajectory. STD conveniently provides pre-grouped trajectories (trails) by applying a time interval-based grouping: for each user, check-ins that occur within a time interval of $\delta_{\tau}=8$ hours are grouped into the same trajectory.
@@
To obtain city-specific datasets, we matched trajectories to the target cities. For each city, we obtain geographic boundaries from OpenStreetMap and retrieve its GeoJSON file via the Overpass API. The GeoJSON file contains a polygon defining the city’s boundary in latitude and longitude. Using this boundary, we filter check-ins by comparing the latitude/longitude of each POI’s administrative region and retain only those that are within the city’s polygon. This ensures that all retained trajectories are spatially grounded within the designated city.
@@
We split trajectories into training, validation, and test sets in a ratio of 7:1:2, following [7]. We ensure that all users in the test set appear at least once in the training or validation set, a common practice in prior studies [60, 54].
3.1.2 POI Enrichment via Foursquare OS Places
@@
Since the POIs in STD include their corresponding Foursquare Place IDs, we matched them directly with entries in the Foursquare OS Places dataset using these IDs as the key. This one-to-one ID correspondence allows for a straightforward join operation, enriching each POI with additional metadata such as its precise latitude and longitude, name (e.g., of a restaurant or subway station), and address. However, not all POIs in the Foursquare OS Places dataset include the full metadata, particularly those categorized as private residences, which are excluded due to privacy restrictions.
3.2 Description and Addressing Literature Gaps
@@
Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset is a city-level POI check-in dataset derived from Semantic Trails [29], comprising user check-in trajectories from 12 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, Moscow, New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. It features anonymized POI check-ins enriched with geographical metadata to support spatiotemporal and sequential modeling tasks. City-level statistics, along with comparisons to existing datasets, are presented in Table 2. Fig. 2 shows a world map highlighting the locations of all cities included in the Massive-STEPS dataset. Table 6 shows the available fields in the dataset and provides an example for each field.
@@
Massive-STEPS offers a more comprehensive and diverse representation of urban mobility compared to typical POI check-in datasets. As shown in Table 2, datasets like FSQ-NYC and FSQ-TKY [57] contain fewer than 10,000 candidate POI locations. In contrast, cities in Massive-STEPS cover significantly more POIs: Massive-STEPS New York has over 49,000 POIs, while Massive-STEPS Jakarta exceeds 76,000. Massive-STEPS Istanbul, one of the largest subsets, features a large user base of 23,700, offering a broad range of user behaviors. Although some Massive-STEPS subsets are smaller than their FSQ counterparts (e.g., Tokyo), we attribute this to the strict filtering procedures applied by STD to remove erroneous entries, as explained in Section 2.1. This scale introduces additional computational challenges. Models that rely on dense POI-to-POI adjacency matrices, for instance, may require more efficient implementations to avoid excessive memory consumption.
@@
Beyond scale, Massive-STEPS addresses the oversaturation of FSQ-NYC and FSQ-TKY in POI recommendation research. Notably, Massive-STEPS includes low-resource and previously underexplored cities in human mobility studies, such as Petaling Jaya and Kuwait City, both of which are among the cities with the highest number of check-ins from STD. This broader coverage opens new research opportunities for studying location-based behaviors across diverse cultural and geographic contexts. Furthermore, since Massive-STEPS is based on STD, it benefits from the carefully filtered, high-quality check-ins and a longer, more recent timespan. These characteristics make Massive-STEPS a more relevant and reliable resource for modeling human mobility patterns.
Massive-STEPS is designed to be easily extended to other geographical regions. Since the data processing code is open-source and fully reproducible, adding a new city only requires its geographic boundaries from OpenStreetMap. Moreover, Massive-STEPS is scalable to higher levels of geographic granularity, enabling the creation of provincial, state, and country-level POI check-in datasets, which support collective mobility studies at broader geographic scales.
4 Benchmark
@@
Model Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
FPMC 0.000 0.026 0.029 0.021 0.062 0.059 0.032 0.026 0.030 0.084 0.075 0.176
RNN 0.085 0.077 0.049 0.087 0.059 0.075 0.061 0.064 0.097 0.055 0.080 0.133

### v3

@@ 3.1. Creation Process
Massive-STEPS is derived from Semantic Trails [^1] Semantic Trails is originally released with a CC0 1.0 Universal license. (Monti et al., 2018), incorporating check-ins from both the 2013 and 2018 subsets. We utilize two additional components from STD: (1) the cities metadata file, which provides the latitude and longitude of administrative regions (e.g., towns, suburbs) along with their corresponding country codes obtained from GeoNames; and (2) the POI category mapping, which links each Foursquare Category ID to its descriptive name (e.g., ”Restaurant”). Based on this metadata, each POI is thus associated with several attributes: Foursquare Place ID, Foursquare Category ID, category name, latitude/longitude of the administrative region, the administrative region name, and the country code. For anonymization and privacy purposes and model training, we apply ordinal encoding to the Place IDs and Category IDs, assigning each a unique integer index.
@@
Most POI trajectory models operate on sequences of check-ins, commonly referred to as trajectories. The model is tasked with predicting the next POIs a user is likely to visit, given the current trajectory. STD conveniently provides pre-grouped trajectories (trails) by applying a time interval-based grouping: for each user, check-ins that occur within a time interval of $\delta_{\tau}=8$ hours are grouped into the same trajectory.
@@
To obtain city-specific datasets, we matched trajectories to the target cities. For each city, we obtain geographic boundaries from OpenStreetMap and retrieve its GeoJSON file via the Overpass API. The GeoJSON file contains a polygon defining the city’s boundary in latitude and longitude. Using this boundary, we filter check-ins by comparing the latitude/longitude of each POI’s administrative region and retain only those that are within the city’s polygon. This ensures that all retained trajectories are spatially grounded within the designated city. For reproducibility purposes, we also include the GeoJSON files with our dataset’s release.
@@
We split trajectories into training, validation, and test sets in a ratio of 7:1:2, following (Feng et al., 2025). We ensure that all users in the test set appear at least once in the training or validation set, following prior studies (Yang et al., 2022c; Yan et al., 2023).
Table 3. Statistics of 15 Massive-STEPS subsets, including the number of users, trajectories, POIs, check-ins, and train, validation, and test trajectory counts. $\mu_{\text{len}}$ denotes the mean number of check-ins per trajectory, and $\mu_{\text{int}}$ denotes the mean time interval between check-ins (in hours). For comparison, we also include statistics from existing Foursquare- and Gowalla-based datasets. We report the version used in (Yan et al., 2023).
City Users Trajectories POIs Check-ins #train #val #test $\mu_{\text{len}}$ $\mu_{\text{int}}$
NYC and Tokyo Check-in Dataset (Yang et al., 2014)
New York 1,048 14,130 4,981 103,941 72,206 1,400 1,347 7.55 7.27
Tokyo 2,282 65,499 7,833 405,000 274,597 6,868 7,038 6.32 5.47
Gowalla (Cho et al., 2011; Yuan et al., 2013)
California 3,957 45,123 9,690 238,369 154,253 3,529 2,780 5.24 8.37
Massive-STEPS
Bandung 3,377 55,333 29,026 161,284 38,732 5,534 11,067 2.91 3.17
Beijing 56 573 1,127 1,470 400 58 115 2.57 3.10
Istanbul 23,700 216,411 53,812 544,471 151,487 21,641 43,283 2.52 4.36
Jakarta 8,336 137,396 76,116 412,100 96,176 13,740 27,480 3.00 2.81
Kuwait City 9,628 91,658 17,180 232,706 64,160 9,166 18,332 2.54 5.31
Melbourne 646 7,864 7,699 22,050 5,504 787 1,573 2.80 3.27
Moscow 3,993 39,485 17,822 105,620 27,639 3,949 7,897 2.67 3.36
New York 6,929 92,041 49,218 272,368 64,428 9,204 18,409 2.96 3.16
Palembang 267 4,699 4,343 14,467 3,289 470 940 3.08 3.17
Petaling Jaya 14,308 180,410 60,158 506,430 126,287 18,041 36,082 2.81 2.96
São Paulo 5,822 89,689 38,377 256,824 62,782 8,969 17,938 2.86 3.54
Shanghai 296 3,636 4,462 10,491 2,544 364 728 2.89 3.02
Sydney 740 10,148 8,986 29,900 7,103 1,015 2,030 2.95 3.33
Tangerang 1,437 15,984 12,956 45,521 11,188 1,599 3,197 2.85 3.24
Tokyo 764 5,482 4,725 13,839 3,836 549 1,097 2.52 5.16
POI Enrichment via Foursquare OS Places
@@
Since the POIs in STD include their corresponding Foursquare Place IDs, we matched them directly with entries in the Foursquare OS Places dataset using these IDs as the key. This one-to-one ID correspondence allows for a straightforward join operation, enriching each POI with additional metadata such as its precise latitude and longitude, name (e.g., of a restaurant or subway station), and address [^2] Foursquare OS Places is originally released with an Apache 2.0 license.. However, not all POIs in the Foursquare OS Places dataset include the full metadata, particularly those categorized as private residences, which are expectedly excluded due to privacy restrictions.
3.2. Description and Addressing Literature Gaps
@@
Massive-STEPS is a city-level POI check-in dataset consisting of real-world, user check-in trajectories from 15 cities: Bandung, Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, Moscow, New York, Palembang, Petaling Jaya, São Paulo, Shanghai, Sydney, Tangerang, and Tokyo. It features anonymized POI check-ins enriched with geographical metadata to support spatiotemporal and sequential modeling tasks. City-level statistics, along with comparisons to existing datasets, are presented in Table 3. Table 9 shows the available fields in the dataset and provides an example for each field. Fig. 8 shows a world map highlighting the cities included in the dataset.
@@
Massive-STEPS offers a more comprehensive representation of urban mobility compared to typical POI check-in datasets. As shown in Table 3, datasets like FSQ-NYC and FSQ-TKY (Yang et al., 2014) contain fewer than 10,000 candidate POI locations. In contrast, cities in Massive-STEPS cover significantly more POIs: Massive-STEPS New York has over 49,000 POIs, while Massive-STEPS Jakarta exceeds 76,000. Massive-STEPS Istanbul, one of the largest subsets, features a large user base of 23,700, offering a broad range of user behaviors. Although some Massive-STEPS subsets are smaller than their FSQ counterparts (e.g., Tokyo), we attribute this to the strict filtering procedures applied by STD to remove erroneous entries, as explained in Section 2. Analysis of the dataset reveals a long-tailed distribution in both trajectory length (Fig. 3) and user activity (Fig. 4). Most users exhibit cold-start behavior, contributing only a few short trajectories per outing. This prevalence of short and sparse trajectories underscores the necessity for models capable of learning and generalizing effectively from limited sequential data.
Massive-STEPS provides temporal coverage across two non-consecutive periods, spanning 2012-2013 and 2017-2018 (24 months in total), which enables longitudinal analysis of urban mobility and POI visitation patterns. This temporal span allows us to evaluate how POI recommendation models perform across different time periods (see Section 4.2.3). POIs are inherently dynamic, with substantial closure rates observed in major cities such as New York, Melbourne, and Sydney (see Table 2), indicating that the urban POI landscape is far from static and underscoring the importance of multi-period datasets like ours. Furthermore, we observe clear distributional shifts in the most visited POI categories between the two periods, as illustrated in Fig. 2, indicating that visitation behaviors can change substantially even within the same city
@@
Beyond its scale, Massive-STEPS addresses the oversaturation of FSQ-NYC and FSQ-TKY in POI trajectory modeling research. Notably, Massive-STEPS includes low-resource and previously underexplored cities in human mobility studies, such as Petaling Jaya and Kuwait City, both of which are among the cities with the highest number of check-ins from STD. This broader coverage opens new research opportunities for studying location-based behaviors across diverse cultural and geographic contexts. Furthermore, since Massive-STEPS is based on STD, it benefits from the carefully filtered, high-quality check-ins and a longer, more recent timespan. These characteristics make Massive-STEPS a more relevant and reliable resource for modeling human mobility patterns.
@@
Model Bandung Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Palembang Petaling Jaya São Paulo Shanghai Sydney Tangerang Tokyo
A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5
FPMC 0.048 0.083 0.000 0.009 0.026 0.050 0.029 0.058 0.021 0.054 0.062 0.107 0.059 0.094 0.032 0.061 0.102 0.136 0.026 0.057 0.030 0.055 0.084 0.120 0.075 0.131 0.104 0.166 0.176 0.239
RNN 0.062 0.099 0.085 0.134 0.077 0.130 0.049 0.083 0.087 0.146 0.059 0.083 0.075 0.122 0.061 0.092 0.049 0.085 0.064 0.107 0.097 0.147 0.055 0.090 0.080 0.125 0.087 0.135 0.133 0.197
[…]
Flashback 0.103 0.155 0.278 0.359 0.137 0.200 0.113 0.166 0.150 0.227 0.079 0.115 0.121 0.180 0.086 0.123 0.135 0.198 0.072 0.119 0.159 0.222 0.096 0.132 0.127 0.174 0.185 0.262 0.170 0.237
Graph-Flashback 0.091 0.146 0.304 0.373 0.137 0.202 0.072 0.121 0.115 0.181 0.093 0.143 0.125 0.188 0.059 0.093 0.148 0.208 0.072 0.119 0.096 0.156 0.111 0.151 0.135 0.188 0.200 0.292 0.180 0.259
LoTNext 0.076 0.125 0.243 0.320 0.105 0.168 0.044 0.078 0.132 0.209 0.085 0.133 0.089 0.145 0.061 0.099 0.140 0.208 0.055 0.097 0.105 0.166 0.093 0.139 0.123 0.176 0.126 0.190 0.191 0.258
[…]
STHGCN 0.219 0.302 0.453 0.552 0.241 0.318 0.197 0.270 0.225 0.314 0.168 0.247 0.223 0.308 0.146 0.207 0.246 0.341 0.174 0.241 0.250 0.344 0.193 0.264 0.227 0.307 0.293 0.400 0.250 0.350
UniMove 0.007 0.033 0.036 0.128 0.015 0.038 0.004 0.020 0.023 0.073 0.008 0.037 0.009 0.030 0.004 0.016 0.009 0.035 0.008 0.034 0.002 0.009 0.000 0.029 0.015 0.059 0.001 0.029 0.032 0.072
Massive-STEPS is designed to be easily extended to other geographical regions. Since the data processing code is open-source and fully reproducible, adding a new city only requires its geographic boundaries from OpenStreetMap. Moreover, Massive-STEPS is scalable to higher levels of geographic granularity, enabling the creation of provincial, state, and country-level POI check-in datasets, which support collective mobility studies at broader geographic scales. To ensure full transparency and facilitate further research, all benchmarking scripts and evaluation code used in our experiments are also publicly released.
Figure 3. Distribution of trail lengths, showing a long-tailed pattern with most trajectories consisting of a few check-ins.
Distribution of trail lengths, showing a long-tailed pattern with most trajectories consisting of a few check-ins.
Figure 4. Distribution of user activity based on the number of trajectories per user, indicating a cold-start-heavy dataset.
Distribution of user activity based on the number of trajectories per user, indicating a cold-start-heavy dataset.
4. Benchmark Tasks
4.1. POI Recommendation
This benchmark focuses on POI recommendation, where the goal is to predict a user’s next visit based on their previous check-ins. The input is a trajectory of visited places, and the model is expected to suggest a set of $K$ POIs the user might visit next. It is a supervised task, trained on all available historical trajectories to learn personalized movement patterns. Appendix C provides details on problem formulation, hyperparameters, experimental setups, and full evaluation results.
4.1.1. Experimental Setup
We adopted the predefined trajectories from the original STD, where check-ins are grouped into sequences based on fixed time intervals (see Section 3.1). All input features are numerically encoded, enabling straightforward use across experiments. Models typically use four feature types: (1) social: user ID; (2) spatial: POI ID and geographic coordinates; (3) temporal: check-in timestamp; and (4) categorical: POI category. For POIs lacking exact geographic coordinates due to privacy (see Section 3.1), we defaulted to the coordinates of their administrative region as a spatial proxy. We evaluated four kinds of architectures: (1) Markov-based methods: FPMC (Rendle et al., 2010), (2) classical deep learning models: RNN (Wang et al., 2021a), LSTPM (Sun et al., 2020), DeepMove (Feng et al., 2018), Flashback (Yang et al., 2021), Graph-Flashback (Rao et al., 2022), LoTNext (Xu et al., 2024), (3) Transformer-based graph neural networks: GETNext (Yang et al., 2022c), STHGCN, and (4) trajectory foundation model: UniMove (Han et al., 2025b). We employed two commonly used metrics in POI recommender systems: Acc@k, which checks if the true POI appears in the top-k predicted results, and NDCG@k, which measures the ranking quality of the suggested results.
[…]
Method Bandung Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Palembang Petaling Jaya São Paulo Shanghai Sydney Tangerang Tokyo
A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5
[…]
Qwen 2.5 7B 0.060 0.111 0.058 0.218 0.035 0.148 0.105 0.179 0.080 0.155 0.030 0.083 0.090 0.185 0.070 0.131 0.075 0.143 0.030 0.116 0.090 0.188 0.040 0.108 0.035 0.091 0.095 0.196 0.110 0.243
Llama 3.1 8B 0.010 0.055 0.000 0.000 0.020 0.065 0.055 0.104 0.030 0.066 0.010 0.040 0.030 0.068 0.025 0.061 0.005 0.025 0.010 0.050 0.030 0.098 0.005 0.013 0.020 0.053 0.020 0.073 0.005 0.025
[…]
Qwen 2.5 7B 0.055 0.126 0.038 0.237 0.040 0.141 0.065 0.161 0.050 0.140 0.040 0.100 0.080 0.176 0.050 0.116 0.050 0.135 0.045 0.111 0.095 0.198 0.045 0.103 0.045 0.109 0.100 0.215 0.120 0.257
Llama 3.1 8B 0.045 0.131 0.077 0.221 0.040 0.137 0.045 0.126 0.060 0.137 0.040 0.101 0.080 0.183 0.055 0.111 0.070 0.154 0.030 0.123 0.030 0.159 0.060 0.116 0.040 0.110 0.080 0.173 0.110 0.269
Gemma 2 9B 0.065 0.130 0.096 0.217 0.045 0.141 0.105 0.180 0.070 0.153 0.050 0.100 0.080 0.194 0.075 0.129 0.060 0.150 0.065 0.126 0.075 0.192 0.050 0.112 0.045 0.103 0.100 0.227 0.110 0.263
[…]
Llama 3.1 8B 0.030 0.033 0.058 0.100 0.015 0.036 0.015 0.021 0.010 0.023 0.040 0.123 0.005 0.031 0.035 0.084 0.010 0.013 0.040 0.049 0.045 0.045 0.020 0.030 0.055 0.141 0.000 0.003 0.030 0.046
[…]
Figure 5. Spearman correlation between category entropy and mean score of POI recommendation models. Comparison against other city features are provided in Appendix D.
Spearman correlation between category entropy and mean score of POI recommendation models.
4.1.2. Main Results
As shown in Table 4, STHGCN achieves the highest average Acc@1 and NDCG@5 across all cities, followed closely by GETNext, demonstrating the effectiveness of GNNs compared to classical recurrent models. The top model attained a mean Acc@1 of 23.4%, comparable to previous studies on similarly sized datasets (Feng et al., 2025). Interestingly, newer variants of Flashback, such as Graph-Flashback and LoTNext, do not always provide a significant improvement, contrary to their original claims. We note that in their respective studies, these methods were evaluated only on two datasets (Gowalla and Foursquare); our results further highlight the importance of benchmarking breadth to ensure that improvements generalize across datasets and diverse urban environments. Similarly, pre-training UniMove from scratch [^3] UniMove pre-trained weights are not publicly available. struggled to surpass older baselines. We attribute this to the high number of cold-start trajectories (see Fig. 3), as the method fails to generalize to the extremely short input sequences typical of real-world data. In these cases, next-token prediction loss struggles to provide meaningful training signal.
4.1.3. Correlating Urban Features with Model Scores
We also examined the impact of urban features on POI recommendation accuracy by computing Spearman correlations between city features and model performance. As shown in Fig. 11, we found that category entropy, based on Shannon entropy, shows a strong negative correlation with predictive accuracy (r=0.646,p<0.05r=-0.646,p<0.05). Our finding suggests that cities with more evenly distributed POI categories tend to be harder to predict. Conversely, cities dominated by a few specific POI categories, such as those with a significant density of Cafés, Restaurants, etc., tend to be easier to model and predict. Intuitively, these specialized urban environments often facilitate more regular and predictable travel behaviors, as users frequently visit these common venue types as part of their daily routines. Further details regarding the relationship between category entropy and model accuracies are provided in Appendix D.
4.2. Zero-shot POI Recommendation
This benchmark focuses on zero-shot POI recommendation via LLMs, where the goal is to predict a user’s next visit based on their previous check-ins (similar to its supervised counterpart) without additional model fine-tuning. The input is a user trajectory transformed into a textual prompt, and the model ranks a set of $K$ candidate POIs to identify the next likely destination. Appendix E provides details on problem formulation, prompts, experimental setups, and full evaluation results.
Figure 6. Acc@1 of supervised (solid lines) and LLM-Move (dashed lines) models across 15 cities.
Acc@1 of supervised (solid lines) and LLM-Move (dashed lines) models across 15 cities.
4.2.1. Experimental Setup
For zero-shot recommendation, trajectories are converted into textual prompts (Xue et al., 2022; Xue and Salim, 2024). We adapted the prompt templates from (Feng et al., 2025), which implemented the three LLM methods evaluated in this study: LLM-Mob (Wang et al., 2023c), LLM-ZS (Beneduce et al., 2024), and LLM-Move (Feng et al., 2024a). Since LLMs can leverage contextual information, features do not need numerical encoding; we used each check-in’s timestamp, POI category name, and POI ID. For a robust evaluation, we tested each method on four LLMs: one closed-source API (Gemini 2.0 Flash (Team and et al., 2024a)) and three open-source instruction-tuned models: Qwen 2.5 7B (Team, 2024), Llama 3.1 8B (Grattafiori et al., 2024), and Gemma 2 9B (Team and et al., 2024b). We used the same metrics as in the supervised setting.
4.2.2. Main Results
As shown in Table 5, LLM-Move (Feng et al., 2024a) outperformed the other two methods due to its prompt design which provides candidate POIs rather than relying solely on historical or contextual trajectories, unlike LLM-Mob and LLM-ZS. Importantly, LLM-Move selects the top $N$ closest locations from a user’s last check-in, using a distance-based heuristic echoing TimeGeo (Jiang et al., 2016). Across LLMs, Gemini 2.0 Flash achieved the highest accuracy across all prompting strategies, with Qwen 2.5 7B and Gemma 2 9B as strong open-source alternatives. Notably, as shown in Fig. 6, these zero-shot methods matched or exceeded supervised baselines in several cities (e.g., Jakarta, Kuwait City, Moscow), demonstrating their effectiveness without fine-tuning.
[…]
Time Period Model Bandung Beijing Istanbul Jakarta KC Melbourne Moscow NY Palembang PJ SP Shanghai Sydney Tangerang Tokyo
[…]
Llama 3.1 8B 0.031 0.041 0.007 0.010 0.000 0.039 0.005 0.032 0.010 0.014 0.048 0.006 0.064 0.000 N/A